OpenStudy (anonymous):

Pleaseee help! :OO A science experiment begins with a metal at -100° Celsius. The following function describes the temperature change per minute: f(x) = 89x - 100°. How will the graph of this function change if the metal is at 25° at the start of the experiment? The line will shift vertically up by 125°. The line will shift vertically up by 25°. The line will shift vertically down by 125°. The line will shift vertically down by 25°.

OpenStudy (anonymous):

Please help... :((( I know it's probably a really easy question but i need help bad im really dumb at this stuff.

OpenStudy (jack1):

y = mx + c equation for a straight line m = gradient c = y intercept (ie when x = 0) so your equation is y = 89x - 100 m = 89 c = (-100) so neg 100 degrees is your initial starting point if you make +25 degrees as your starting point: new eqn will be y = mx + c y = 89x + 25 gradient wont change, c value will on a graph it would look something like below (really rough drawing) 

OpenStudy (jack1):

as you can see, the new line has shifted UPWARDS and its moved by : 

OpenStudy (jack1):

is that all making sense @barbieeebaby ?

OpenStudy (anonymous):

oh my god you're amazing thank-you so much! I really appreciate your straight forward answer. and yeah, it all really helped! I'll try your method on the other problems! thanks sooo much again! :))

OpenStudy (jack1):

nah all good hey, happy to help ;D good luck with the rest

OpenStudy (anonymous):

Okay! Thanks! :DDD